A few years ago a total of 2684 thousand people lived in the metropolitan areas of Las Vegas, Nevada, and Sacramento, California. Sacramento had 282 thousand more residents than Las Vegas. What was the population of each metropolitan area?

Answer by checkley79(3341)

X=POPULATION OF LAS VEGAS.
X+282,000 =POPULATION OF SACRAMENTO.
2,684,000=X+(X+282,000)
2,684,000=2X+282,000
2,684,000-282,000=2X
2,402,000=2X
2,402,000/2=X
1,201,000=X THE POPULATION OF LAS VEGAS.
1,201,000+282,000=1,483,000 THE POPULATION OF SACRAMENTO.
PROOF:
2,684,000=1,201,000+1,483,000
2,684,000=2,684,000